You are the Financial Manager of the Trading Division of Eager Foreign Exchange Division. Your Latin American Division in Houston is providing you the following quotations: (You may either buy or sell at the stated rates). Bank of New York: Peso 10.00/$ Bank of Cancun: Sol 0.33/ Peso Bank of Lima: $0.33/ Sol Assume you have an initial wealth of $1,000,000 U.S. dollars in U.S.A. Is Triangular Arbitrage Possible with Latin American Currencies? If so, compute Eager’s profits?
Answer #1

Initial wealth is $1,000,000.

This is converted into Peso at the Peso/$ rate. Peso received = $1,000,000 * 10 = 10,000,000 Peso.

10,000,000 Peso is converted into Sol at the Sol/$ rate. Sol received = 10,000,000 * 0.33 = 3,300,000 Sol.

3,300,000 Sol is converted into $ at the $/Sol rate. $ received = 3,300,000 * 0.33 = $1,089,000.

Profit = $1,089,000 - $1,000,000 = $89,000.